# Generalised 4d Partition Functions and Modular Differential Equations

### v1

@@ 1 Introduction and Summary
Rather surprisingly, the Schur index of a 4d ${\cal N}=2$ SCFT can also be computed by an infrared formula as the trace of the inverse of the Kontsevich-Soibelmann monodromy operator [8] (also see [9, 10, 11, 12]). Within this framework, it has been observed in examples that traces of higher powers, say $k$, of the monodromy operator yield $q$-series that can be identified with vacuum characters of certain vertex algebras ${\cal V}^{(k)}$ [13, 14]. While k=1k=-1 recovers the non-unitary VOA arising in the 4d/2d correspondence, the monodromy traces for $k>0$ encountered in many examples match the characters of unitary VOAs, including rational and logarithmic 2d conformal field theories.
@@ 2.2.2 Schur indices for gauge theories
$\begin{split}{\cal I}(q)=\frac{(-i)^{{\rm rk}\,G-{\rm dim}\,G}\,\eta(\tau)^{3\,{\rm rk}\,G-{\rm dim}\,G}}{|W|}\oint\prod_{j=1}^{{\rm rk}\,\mathfrak{g}}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\prod_{\alpha\neq 0}\vartheta_{1}(\alpha(\mathbf{u}),\tau)\prod_{l}\prod_{\rho\in\mathcal{R}_{l}}\left(\frac{\eta(\tau)}{\vartheta_{4}(\rho(\mathbf{u}),\tau)}\right)^{N_{l}}~,\end{split}$ (2.14)
where $\alpha$ denotes the non-zero roots, $\rho$ denotes the weights of the representation $\mathcal{R}_{l}$, and we define
@@
Schematically, the contribution of the vector multiplet and the Haar measure in Eq. (2.3) combine to produce the overall power of the Dedekind eta function $\eta(\tau)$ and the product over $\vartheta_{1}(\alpha(\mathbf{u}),\tau)$ in Eq. (2.14). On the other hand, the hypermultiplet contribution can be simplifed into the ratio of $\eta(\tau)$ and $\vartheta_{4}(\rho(\mathbf{u}),\tau)$.
@@ 2.3 Generalised Schur partition function
$\begin{split}\widehat{{\cal Z}}_{G}(q;\alpha)&=\frac{1}{N(\alpha)\,|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\,\Delta(\mathbf{z})\,{\rm PE}\big[i_{V}(q,\mathbf{z};\alpha)+\sum_{l}N_{l}\,i_{H}^{\mathcal{R}_{l}}(q,\mathbf{z};\alpha)\big]\\ &=\frac{(q;q)^{2\,{\rm rk}\,G}}{N(\alpha)\,|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\,\Delta(\mathbf{z})^{\alpha}\left(\frac{\prod_{\alpha\neq 0}(q\,\mathbf{z}^{\alpha};q)^{2}}{\prod_{l}\prod_{\rho\in\mathcal{R}_{l}}(\mathbf{z}^{\rho}\sqrt{q};q)^{N_{l}}}\right)^{\alpha}~,\end{split}$ (2.21)
where $N(\alpha)$ is a constant prefactor introduced by hand such that the leading order coefficient of this function is 1, and is defined to be [^6] This is a trivial generalisation of the normalisation considered in [15] to a general gauge group $G$.
@@
$\begin{split}&{\cal Z}_{G}(q;\alpha)\\ &\quad=C(\alpha)\,\frac{\eta(\tau)^{2\,{\rm rk}\,G}}{|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\bigg(\eta(\tau)^{{\rm rk}\,G-{\rm dim}\,G}\prod_{\alpha\neq 0}\vartheta_{1}(\alpha(\mathbf{u}),\tau)\prod_{l}\prod_{\rho\in\mathcal{R}_{l}}\frac{\eta(\tau)^{N_{l}}}{\vartheta_{4}^{N_{l}}(\rho(\mathbf{u}),\tau)}\bigg)^{\alpha}~,\end{split}$ (2.23)
@@ 2.5 Contour integral representation for RCFT characters
The constants ${\cal N}_{A}$ are a set of normalisations that are worked out in detail in [18]. The contour integrals are to be performed in the order $t_{1},t_{2},\cdots t_{n-1}$ and the subscripts ${\cal C}_{A}$ means the contours for $t_{1},\cdots t_{A}$ run from 0 to $\lambda$, while those for $t_{A+1},\cdots t_{n-1}$ run from 1 to $\infty$. In particular, this means that for A=n1A=n-1 (the maximum allowed value), all contours run from 0 to $\lambda$.
@@ 3 $SU(2)$ SCFT with $4$ fundamental hypers
and equating it to 6-6, we land on $k=-\frac{4}{3}$. Thus one can reinterpret the character in Eq. (3.26) as either the second/non-identity character of $SU(3)_{1}$, or the first/identity character of $V_{-\frac{4}{3}}(SU(2))$. In the latter form, it has been shown to correspond to the Schur index of the $(A_{1},A_{3})$ Argyres-Douglas theory [8]. It should be kept in mind that this type manipulation of exchanging unitary and non-unitary characters does not necessarily lead to a meaningful interpretation in terms of RCFT/VOAs in every case.
@@ 4 $USp(2N)$ SCFT with $2N+2$ fundamental hypers
$\prod_{\alpha\neq 0}\vartheta_{1}(\alpha(\mathbf{u}),\tau)=\prod_{i=1}^{N}\vartheta_{1}(\pm 2u_{i})\prod_{1\leq i<j\leq N}\vartheta_{1}(\pm u_{i}\pm u_{j})~.$ (4.2)
@@ 4.1 First non-trivial case: $N=2$
$\mu_{2}=\frac{1}{2},\quad\mu_{3}=\frac{1}{216}(1-8\alpha)(1-2\alpha)(1+10\alpha)$ (4.25)
@@
Another interesting case is the non-unitary minimal model ${\cal M}(2,7)$ with central charge $c=-\frac{68}{7}$ and $(h_{1},h_{2})=(-\frac{2}{7},-\frac{3}{7})$. In this form, these numbers do not fit our formula Eq. (4.24). However one can change the “presentation” by treating the third character as the identity (this procedure is discussed in [50] for the three-character case). Then the effective “central charge” changes to $c-24h_{2}=\frac{4}{7}$. Now the conformal dimensions of the other characters become $\frac{1}{7},\frac{3}{7}$ and we have a perfect match for $\alpha=\frac{1}{7}$. Note however that this change of presentation, despite making $c$ and all the $h_{A}$ positive, does not render the theory unitary. The vacuum character of this non-unitary minimal model is the associated VOA of the $(A_{1},A_{4})$ Argyres–Douglas theory [8].
@@ 4.2 General NN
which agrees with $12c_{4d}$ as quoted in [22].
@@
Thus we have reproduced all four cases in Table III of [22] corresponding to the values $\alpha=\frac{1}{2},1,\frac{1}{2N+3},\frac{1}{2N+1}$ of the parameter of the generalised Schur partition function.
@@ 4.3 Two-parameter generalised Schur partition function
$\begin{split}&{\cal Z}_{USp(2N)}(q;\alpha,\beta)=\\ &\quad\times\;\frac{\eta^{2N}}{|W|\,N(\alpha)}\oint\prod_{j=1}^{N}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\prod_{i=1}^{N}\left(\eta^{6}\,\frac{\vartheta_{1}^{2}(2u_{i})}{\vartheta_{4}^{8}(u_{i})}\right)^{\alpha}\prod_{1\leq i<j\leq N}\left(\eta^{4}\,\frac{\vartheta_{1}^{2}(u_{i}\pm u_{j})}{\vartheta_{4}^{4}(u_{i})\vartheta_{4}^{4}(u_{j})}\right)^{\beta}~.\end{split}$ (4.46)
@@ 5 Discussion and open questions
We note here that the appearance of characters of unitary RCFTs in ${\cal Z}_{USp(2N)}(q;\alpha)$ for special values of $\alpha$, as discussed in the main body of this work, is intriguing. Although we demostrate this by showing that they arise as solutions to the same one-parameter sub-family of an MLDE, a direct 4d interpretation of these SCFT/unitary RCFT connections remains to be understood. [^18] A promising direction, based on the discussion in Subsection 5.1 might be to consider twisted circle compactifications based on and explored in the following series of works [64, 65, 66, 67, 68, 69, 70, 71].
@@ 5.1 Relation to BPS monodromy traces and MLDE
For any given 4d ${\cal N}=2$ superconformal field theory, the following trace of $k$ powers of the associated BPS monodromy matrix:
@@
solves a modular linear differential equation of fixed order and Wronksian index, for all $k$. Here, $r$ is the rank of the SCFT and
@@
It is important to remember here that for large $|k|$, ${\rm Tr}\,M^{k}$ diverges and this conjecture is subject to an appropriate regularisation. For k=1k=-1, this reduces to the statement that the Schur index, which itself is conjecturally equal to ${\cal I}_{-1}(q)$, solves an MLDE – this has been checked extensively. Thus according to Conjecture 1, the MLDE solved by traces of higher powers of the BPS monodromy operator has the same order and Wronskian index as the one solved by the Schur index.
@@ A.1 Jacobi theta functions
$(z;q)\equiv(z;q)_{\infty}:=\prod_{i=0}^{\infty}(1-zq^{n})$ (A.14)

### v2

@@ 1 Introduction and Summary
Rather surprisingly, the Schur index of a 4d ${\cal N}=2$ SCFT can also be computed by an infrared formula as the trace of the inverse of the Kontsevich-Soibelmann monodromy operator [18, 30] (also see [23, 52, 17, 29, 28, 19]). Within this framework, it has been observed in examples that traces of higher powers, say $k$, of the monodromy operator yield $q$-series that can be identified with vacuum characters of certain vertex algebras ${\cal V}^{(k)}$ [24, 56]. While k=1k=-1 recovers the non-unitary VOA arising in the 4d/2d correspondence, the monodromy traces for $k>0$ encountered in many examples match the characters of unitary VOAs, including rational and logarithmic 2d conformal field theories.
@@ 2.2.2 Schur indices for gauge theories
$\begin{split}{\cal I}(q)=\frac{(-i)^{{\rm rk}\,G-{\rm dim}\,G}\,\eta(\tau)^{3\,{\rm rk}\,G-{\rm dim}\,G}}{|W|}\oint\prod_{j=1}^{{\rm rk}\,\mathfrak{g}}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\prod_{\bm{\alpha}\neq 0}\vartheta_{1}(\bm{\alpha}\cdot\mathbf{u},\tau)\prod_{l}\prod_{{\bm{\rho}}\in\mathcal{R}_{l}}\left(\frac{\eta(\tau)}{\vartheta_{4}(\bm{\rho}\cdot\mathbf{u}),\tau)}\right)^{N_{l}}~,\end{split}$ (2.14)
where $\bm{\alpha}$ denotes the non-zero roots, $\bm{\rho}$ denotes the weights of the representation $\mathcal{R}_{l}$, and we define
@@
Schematically, the contribution of the vector multiplet and the Haar measure in Eq. (2.3) combine to produce the overall power of the Dedekind eta function $\eta(\tau)$ and the product over $\vartheta_{1}(\bm{\alpha}\cdot\mathbf{u},\tau)$ in Eq. (2.14). On the other hand, the hypermultiplet contribution can be simplifed into the ratio of $\eta(\tau)$ and $\vartheta_{4}(\bm{\rho}\cdot\mathbf{u},\tau)$.
@@ 2.3 Generalised Schur partition function
$\begin{split}\widehat{{\cal Z}}_{G}(q;\alpha)&=\frac{1}{N(\alpha)\,|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\,\Delta(\mathbf{z})\,{\rm PE}\big[i_{V}(q,\mathbf{z};\alpha)+\sum_{l}N_{l}\,i_{H}^{\mathcal{R}_{l}}(q,\mathbf{z};\alpha)\big]\\ &=\frac{(q;q)^{2\,{\rm rk}\,G}}{N(\alpha)\,|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\,\Delta(\mathbf{z})^{\alpha}\left(\frac{\prod_{\bm{\alpha}\neq 0}(q\,e^{2\pi i\,\bm{\alpha}\cdot\mathbf{u}};q)^{2}}{\prod_{l}\prod_{{\bm{\rho}}\in\mathcal{R}_{l}}(e^{2\pi i\,\bm{\rho}\cdot\mathbf{u}}\sqrt{q};q)^{N_{l}}}\right)^{\alpha}~,\end{split}$ (2.21)
where recall that $z_{j}=e^{2\pi i\,u_{j}}$, while $N(\alpha)$ is a constant prefactor introduced by hand such that the leading order coefficient of this function is 1, and is defined to be [^6] This is a trivial generalisation of the normalisation considered in [34] to a general gauge group $G$.
@@
$\begin{split}&{\cal Z}_{G}(q;\alpha)\\ &\quad=C(\alpha)\,\frac{\eta(\tau)^{2\,{\rm rk}\,G}}{|W|}\oint\prod_{j=1}^{{\rm rk}\,G}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\bigg(\eta(\tau)^{{\rm rk}\,G-{\rm dim}\,G}\prod_{\bm{\alpha}\neq 0}\vartheta_{1}(\bm{\alpha}\cdot\mathbf{u},\tau)\prod_{l}\prod_{\rho\in\mathcal{R}_{l}}\frac{\eta(\tau)^{N_{l}}}{\vartheta_{4}^{N_{l}}(\bm{\rho}\cdot\mathbf{u}),\tau)}\bigg)^{\alpha}~,\end{split}$ (2.23)
@@ 2.5 Contour integral representation for RCFT characters
The constants ${\cal N}_{A}$ are a set of normalisations that are worked out in detail in [64]. The contour integrals are to be performed in the order $t_{1},t_{2},\cdots t_{n-1}$ and the subscript ${\cal C}_{A}$ means the contours for $t_{1},\cdots t_{A}$ run from 0 to $\lambda$, while those for $t_{A+1},\cdots t_{n-1}$ run from 1 to $\infty$. In particular, this means that for A=n1A=n-1 (the maximum allowed value), all contours run from 0 to $\lambda$.
@@ 3 $SU(2)$ SCFT with $4$ fundamental hypers
and equating it to 6-6, we land on $k=-\frac{4}{3}$. Thus one can reinterpret the character in Eq. (3.26) as either the second/non-identity character of $SU(3)_{1}$, or the first/identity character of $V_{-\frac{4}{3}}(SU(2))$. In the latter form, it has been shown to correspond to the Schur index of the $(A_{1},A_{3})$ Argyres-Douglas theory [18, 30]. It should be kept in mind that this type manipulation of exchanging unitary and non-unitary characters does not necessarily lead to a meaningful interpretation in terms of RCFT/VOAs in every case.
@@ 4 $USp(2N)$ SCFT with $2N+2$ fundamental hypers
$\prod_{\bm{\alpha}\neq 0}\vartheta_{1}(\bm{\alpha}\cdot\mathbf{u},\tau)=\prod_{i=1}^{N}\vartheta_{1}(\pm 2u_{i})\prod_{1\leq i<j\leq N}\vartheta_{1}(\pm u_{i}\pm u_{j})~.$ (4.2)
@@ 4.1 First non-trivial case: $N=2$
$\mu_{2}=\frac{1}{36}-\frac{7\alpha^{2}}{3},\quad\mu_{3}=\frac{1}{216}(1-8\alpha)(1-2\alpha)(1+10\alpha)$ (4.25)
@@
Another interesting case is the non-unitary minimal model ${\cal M}(2,7)$ with central charge $c=-\frac{68}{7}$ and $(h_{1},h_{2})=(-\frac{2}{7},-\frac{3}{7})$. In this form, these numbers do not fit our formula Eq. (4.24). However one can change the “presentation” by treating the third character as the identity (this procedure is discussed in [65] for the three-character case). Then the effective “central charge” changes to $c-24h_{2}=\frac{4}{7}$. Now the conformal dimensions of the other characters become $\frac{1}{7},\frac{3}{7}$ and we have a perfect match for $\alpha=\frac{1}{7}$. Note however that this change of presentation, despite making $c$ and all the $h_{A}$ positive, does not render the theory unitary. The vacuum character of this non-unitary minimal model is the associated VOA of the $(A_{1},A_{4})$ Argyres–Douglas theory [30, 19].
@@ 4.2 General NN
which agrees with $12c_{4d}$ as quoted in [Deb:2025].
@@
Thus we have reproduced all four cases in Table III of [Deb:2025] corresponding to the values $\alpha=\frac{1}{2},1,\frac{1}{2N+3},\frac{1}{2N+1}$ of the parameter of the generalised Schur partition function.
@@ 4.3 Two-parameter generalised Schur partition function
$\begin{split}&{\cal Z}_{USp(2N)}(q;\alpha,\beta)=\\ &\quad\;\frac{\eta^{2N}}{|W|\,N(\alpha)}\oint\prod_{j=1}^{N}\frac{{\rm d}z_{j}}{2\pi iz_{j}}\prod_{i=1}^{N}\left(\eta^{6}\,\frac{\vartheta_{1}^{2}(2u_{i})}{\vartheta_{4}^{8}(u_{i})}\right)^{\alpha}\prod_{1\leq i<j\leq N}\left(\eta^{4}\,\frac{\vartheta_{1}^{2}(u_{i}\pm u_{j})}{\vartheta_{4}^{4}(u_{i})\vartheta_{4}^{4}(u_{j})}\right)^{\beta}~.\end{split}$ (4.46)
@@ 5 Discussion and open questions
We note here that the appearance of characters of unitary RCFTs in ${\cal Z}_{USp(2N)}(q;\alpha)$ for special values of $\alpha$, as discussed in the main body of this work, is intriguing and has already been observed in [16]. Although we demonstrate this by showing that they arise as solutions to the same one-parameter sub-family of an MLDE, a direct 4d interpretation of these SCFT/unitary RCFT connections remains to be understood. [^18] A promising direction, based on the discussion in Subsection 5.1 might be to consider twisted circle compactifications based on and explored in the following series of works [41, 36, 4, 3, 51, 50, 49, 48].
@@ 5.1 Relation to BPS monodromy traces and MLDE
For any given 4d ${\cal N}=2$ superconformal field theory, the following trace of $k$-th powers of the associated BPS monodromy matrix:
@@
solves a modular linear differential equation of fixed order and Wronksian index, for all real (and also complex) $k$. Here, $r$ is the rank of the SCFT and
@@
It is important to remember here that for large enough integers $|k|$, ${\rm Tr}\,M^{k}$ diverges and this conjecture is subject to an appropriate regularisation, particularly so in the case of non-integral $k$. For k=1k=-1, this reduces to the statement that the Schur index, which itself is conjecturally equal to ${\cal I}_{-1}(q)$, solves an MLDE – this has been checked extensively. Thus according to Conjecture 1, the MLDE solved by traces of higher powers of the BPS monodromy operator has the same order and Wronskian index as the one solved by the Schur index.
@@ A.1 Jacobi theta functions
$(z;q)\equiv(z;q)_{\infty}:=\prod_{n=0}^{\infty}(1-zq^{n})$ (A.14)